# Semantic Inference-Based Deep Learning and Modeling for Earth Observation: Cognitive Semantic Augmentation Satellite Networks

### v1

@@ Abstract
Earth Observation (EO) systems support vital tasks such as mapping, disaster monitoring, and resource management, yet face challenges in effectively processing and transmitting vast volumes of EO data, particularly in specialized applications like agriculture and real-time disaster response. Domain-adapted Large Language Models (LLMs) offer a solution by enabling data fusion between mass EO data and semantic EO data. Through domain adaptation, LLMs improve the integration and interpretation of diverse data sources, addressing the challenge of processing specialized datasets in applications like agriculture and disaster response. The fusion of these data types, guided by LLMs, enhances the accuracy and relevance of the transmitted information. This paper presents a comprehensive overview of leveraging semantic inference for deep learning and semantic modeling for advanced EO systems. A novel framework is presented for semantic communication in EO satellite networks, aimed at enhancing data transmission efficiency and system performance through cognitive processing techniques. The proposed system leverages Discrete Task-Oriented Joint Source-Channel Coding (DT-JSCC) and Semantic Data Augmentation (SA) integrate cognitive semantic processing with inter-satellite links, enabling efficient analysis and transmission of multispectral imagery for improved object detection, pattern recognition, and real-time decision-making. Cognitive Semantic Augmentation (CSA) is introduced to enhance a system’s capability to process and transmit semantic information, improving feature prioritization, consistency, and adaptation to changing environments and application needs. The end-to-end architecture is designed for next-generation satellite networks, such as those supporting 6G, demonstrating significant improvements in efficiency and accuracy.
@@ I Introduction
The United Nations has acknowledged Earth observation (EO) as a crucial instrument for accomplishing many Sustainable Development Goals [1]. EO systems acquire and analyze data about the globe for numerous civil applications and critical tasks that significantly impact human existence through the use of satellite networks. These tasks include mapping, weather forecasting, risk management, disaster monitoring, natural resource management, and emergency response. The estimated value of the satellite-based EO market is expected to reach 11.3 billion USD by 2031 [2]. The Internet of Things (IoT) has led to significant advancements in information and communication technology, which has increased demand for profitable value-added services and applications. The emergence of big EO data has been facilitated by the progress in sensors, communication, and storage technology. This dataset has a wealth of valuable information, but it also presents some difficulties owing to its distinctive characteristics, such as one-time observations and intricate dimensions. The exploration of picture retrieval strategies is tailored to address the challenges posed by complicated data [3]. The study emphasizes notable obstacles, including the scarcity and diversity of training data, as well as the current algorithms’ inability to imitate human-like cognition.
@@
Although there is a large amount of EO data available, there is a lack of effective conversion into actionable information for specialized sectors such as agriculture or disaster response. Big data systems that have frequent satellite revisits have the potential to revolutionize real-time monitoring [4]. NewSpace constellations [5] are expanding their capabilities beyond the provision of affordable and frequent data. They focus on addressing particular requirements by using specialized satellites and providing data analysis in addition to raw data, serving a broader spectrum of consumers. Furthermore, monitoring strategies need careful consideration of both geographical and temporal dimensions [6]. In terms of space, this entails establishing the specific region or extent that is of interest. The extent of the phenomena being examined might vary from a single spot to an intricate three-dimensional grid. Resolution, which refers to the level of detail in measurements, and expanse, which indicates the region covered, are both important aspects. An example of researching the influence of topography on soil moisture may include the measurement of moisture levels at many locations along a slope, necessitating a high level of spatial resolution. Consequently, current EO systems are facing difficulties in managing the substantial volume of data being sent and fulfilling the diverse demands for the latency at which the data has to be processed. In order to address these significant challenges, the primary strategies involve two actions: firstly, improving the bandwidth of satellite communication networks that facilitate EO systems, and secondly, reducing data volume by employing on-board processors on the satellites to preprocess the EO data.
@@
Utilizing the inter-satellite linkages (ISLs) of the multi-satellite, multi-orbit connectivity systems has been seen as a potential strategy, as stated in the first direction [7]. The integration of several satellites and orbits, spearheaded by non-geostationary satellites (NGSO), is on the verge of transforming worldwide communication. Unlike conventional geostationary satellites (GSO), these systems provide reduced latency and signal loss as a result of their lower orbits. These benefits include the provision of fast internet access in isolated regions, improved capacity for transmitting data via 5G networks [8], and the possibility of establishing direct, high-capacity connections with consumers. Nevertheless, the smooth integration with current GSO and terrestrial networks requires careful and detailed design due to several constraints. Efficient cohabitation across systems needs solutions for spectrum access and restrictions [9]. An additional challenge lies in the efficient design and management of extensive constellations of NGSO [10]. It is crucial to build user equipment that can successfully establish communication with these constellations. This research examines the communication in multi-satellite, multi-orbit connectivity systems, with a special emphasis on NGSO technology. It investigates the possibilities of this technology while recognizing the obstacles associated with its adoption. The data acquired by EO satellites may undergo processing, storage, and transmission to Earth using SatCom satellites positioned in various orbits. In the context of this framework, a remote sensing satellite (or EO satellite) refers to a spacecraft that collects EO-related data from distant sensors and IoT devices. By using the ISLs, several routing channels may be constructed between the remote-sensing satellites and their respective ground stations. The technique of [11] may effectively handle the problem of overload and backhaul bottlenecks that occur at certain nodes in inter-satellite communication (ISC) networks, resulting in improved network capacity.
@@
In relation to the second aspect, the use of semantic communication (SemCom) principles has been recognized as a favorable strategy. The primary purpose of SemCom is to prioritize semantic accuracy above reducing bit/symbol error rates at the receiver side, which is in complete alignment with the core goals of most EO methods. The first principle of SemCom is the extraction of semantic data, wherein pertinent information is initially obtained, condensed, or broken down from unprocessed data. Afterward, just the information related to meaning is transferred to the central core. The PhiSat-1 project presented in [15] by the European Space Agency (ESA) might be regarded as the first demonstration of this idea. Two 6U CubeSats are deployed in this mission to capture Earth photographs. The data is processed inside the CubeSats to filter out information depending on the cloud coverage of the scene. The second concept revolves around semantic-oriented network management, specifically SemCom. SemCom’s objective is to oversee all network functionalities across various layers (PHY, MAC, Network, Transportation, Application) by taking into account the semantics of the data and the requirements of applications related to EO tasks [16].
Irrespective of the intended message of these bits, the aspiration of beyond Shannon [17] beyond the traditional Shannon paradigm, which emphasizes the reliable reception of individual sent bits. Semantic-information feature learning [18] is the key to addressing the utilization of cognitive techniques employed as a means to direct computational resources towards activities of higher importance. The process of achieving dynamic adaptation in semantic compression involves the utilization of a feature learning module and an attention feature module. These modules enable the source encoder to generate a variable number of symbols and modify the capabilities of the source encoder and channel encoder through the use of cognitive techniques. The evolution of semantic communication [19] is traced back to the early 20th century with its continuous growth into the realm of modern communications beyond 5G and 6G technologies. In light of this, there is a significant need to develop more intelligent and efficient communication protocols that can meet the diverse quality of service (QoS) needs. This must be done while addressing the challenge of limited communication bandwidth and computation. The development of an intelligent communication system is considered essential in both industry and academia. Such a system is not limited to memorizing data flows based on rigorous regulations but also aims to comprehend, analyze, and articulate the fundamental semantics. In order to facilitate the implementation of immediate intelligent services in future communication systems, it is advantageous to distill and communicate only the information that is pertinent to the task at hand and precise semantics. The semantic approach effectively reduces the overall latency of the system.
Figure 1: The system model of semantic inference-based earth observation satellite networks.
@@
Fig. 1 presents a conceptual model for a multi-orbit inter-satellite network in the context of Earth observation, highlighting a layered approach to satellite communication and data processing. At the center, Earth observation satellites, operating in different orbits (LEO, MEO, and GEO), gather raw data through ground sensors distributed across various landscapes, such as agricultural fields and maritime regions. This raw data is transmitted to a ”Ground segment,” which includes data centers and user control interfaces for further processing. Data travels through multiple layers of communication within the network. At the base level, the PHY Layer (Physical Layer) ensures the fundamental transmission of data, followed by the MAC Layer (Medium Access Control) responsible for resource allocation. Moving upwards, the Network Layer manages the interconnectivity across the satellite network. The Transportation Layer handles the movement of data across the system, connecting the ground segments to various orbital satellites.
@@
Above these operational layers lies the Semantic Effectiveness Plane, which provides a high-level oversight and analysis of the semantic data flow. Two processes stand out at either end of this system: Semantic Extraction and Semantic Restoration. Data extracted at the ground level undergoes semantic processing, where meaningful patterns or insights are identified and interpreted. This processed semantic information is later restored into useful applications, ranging from environmental monitoring to disaster management. The model suggests a closed-loop system where raw data from ground sensors is transformed into semantically rich insights, enabling diverse applications through seamless communication between satellite networks and ground infrastructure.
We summarize the paper as follows. Section I discusses semantic image processing, focusing on pixel-level categorization to improve applications like autonomous driving, medical imaging, and Earth observation while addressing challenges with complex data and deep-learning techniques for more accurate analysis. Section II examines semantic communication networks and their integration into 6G systems, emphasizing goal-oriented methods and joint source-channel coding to enhance communication efficiency. Section III explores adapting LLMs for domain-specific tasks to optimize semantic transmission, while Section VI presents the Cognitive Semantic EO system model, integrating DT-JSCC and semantic data augmentation for efficient EO data processing in satellite networks.
II Semantic data processing for Earth observation and 6G communication
This section delves into the concept of semantic segmentation, a technique that assigns precise identifiers to individual pixels in photographs. It concentrates on the application of this methodology to the examination of satellite imagery and other types of EO data. The content explores the application of semantic segmentation in a variety of applications, including trajectory facilitation, information retrieval, and picture compression. We also underscore the current advancements in deep learning for the classification of various EO data types using semantic knowledge.
@@ II Semantic data processing for Earth observation and 6G communication
Assigning a unique name to each pixel in an image is a basic task in computer vision known as semantic image processing. Achieving accurate categorization at the pixel level is crucial in applications like driverless cars, where the capacity to understand lane markings and precisely identify the position of objects is essential. However, the process of semantic segmentation may be complicated by barriers such as hidden objects or congested backgrounds. Notwithstanding these challenges, semantic segmentation offers benefits that go beyond autonomous vehicles. The technology enables the examination of medical images by identifying abnormalities, assists robots in navigating their surroundings, and improves video surveillance by detecting potentially troublesome behavior. Recent progress in deep learning has led to remarkable achievements in achieving precise precision at the pixel level. However, its performance is inferior to that of normal satellite images when dealing with non-standard data such as point clouds and spectrum imaging. Likewise, a significant disparity in performance occurs when dealing with limited data sets. The lack of comprehensively classified, unconventional remote sensing data hinders the development and evaluation of innovative deep-learning methods.
II-1 Computational semantic compression
The study presented in [20] investigates potential resolutions for constraints such as streamlining features, indexing, semantic labeling, and ultimately, approaches for formulating and executing searches. As stated in [21], the primary objective of the study is to evaluate the existing condition of research on extracting information from EO data. The authors of [22] provide a new notion of a semantic EO data cube. It enriches traditional exploratory optimization data cubes by including an extra layer, referred to as a ”nominal” or categorical interpretation, for each observation. Thus, it becomes feasible to use semantic searches that directly relate to the substance of the data and are easily comprehensible to people. By using machine-based reasoning, this approach improves automated analysis and information extraction in comparison to non-semantic cubes. Fundamentally, semantic EO data cubes enable a wider range of users to utilize comprehensive EO data. This study evaluates CORSA [23] and [24], an innovative compression method powered by artificial intelligence that produces useful data representations at the same time. The authors compare traditional methods for detecting water with innovative ways that use the compressed data and representations of CORSA. The results suggest that CORSA compression does not have any detrimental effect on accuracy and allows for the direct generation of subsequent water detection networks based on its representations. These results demonstrate the adaptability of CORSA as a solution for processing and compressing data on-board, successfully addressing the growing challenge of managing EO data. Furthermore, the authors of [25] propose a novel method that uses semantic technology to provide a standardized language for EO data. The methodology emphasizes the smooth integration of data obtained from raster imagery, such as land cover maps, with other geographical information. This enhances the potential for more efficient analysis and use of this crucial EO data.
II-2 Fusion of semantic trajectories
Global Positioning System (GPS) data produces huge amounts of trajectory information. While scholars have largely focused on the preservation and analysis of these trajectories, there is a growing need to understand the underlying importance of the movement. The material includes traditional compression algorithms, the criteria for defining moving objects, approaches for measuring compression efficiency, and potential future applications. For real-time applications, conventional methods for producing substantial movement data are inadequately fast. The research presented in [27] presents a novel approach that combines mobility data with geography and application information to produce semantic trajectories. These improved trajectories not only reveal the destinations of objects but also provide a valuable understanding of the fundamental causes of their motion.
II-3 Advancements in Semantic Comprehension

### v3

# On-Air Deep Learning Integrated Semantic Inference Models for Enhanced Earth Observation Satellite Networks: A Review Beyond Horizons

@@ Abstract
Earth Observation (EO) systems are crucial for cartography, disaster surveillance, and resource administration. Nonetheless, they encounter considerable obstacles in the processing and transmission of extensive data, especially in specialized domains such as precision agriculture and real-time disaster response. Earth observation satellites, outfitted with remote sensing technology, gather data from onboard sensors and IoT-enabled terrestrial objects, delivering important information remotely. Domain-adapted Large Language Models (LLMs) provide a solution by enabling the integration of raw and processed EO data. Through domain adaptation, LLMs improve the assimilation and analysis of many data sources, tackling the intricacies of specialized datasets in agriculture and disaster response. This data synthesis, directed by LLMs, enhances the precision and pertinence of conveyed information. This study provides a thorough examination of using semantic inference and deep learning for sophisticated EO systems. It presents an innovative architecture for semantic communication in EO satellite networks, designed to improve data transmission efficiency using semantic processing methodologies. Recent advancements in onboard processing technologies enable dependable, adaptable, and energy-efficient data management in orbit. These improvements guarantee reliable performance in adverse space circumstances using radiation-hardened and reconfigurable technology. Collectively, these advancements enable next-generation satellite missions with improved processing capabilities, crucial for operational flexibility and real-time decision-making in 6G satellite communication.
@@ I Introduction
Earth Observation (EO) systems collect and analyze global data for many civic applications and essential services that profoundly affect human life via satellite networks. The duties include cartography, meteorological forecasting, risk assessment, disaster surveillance, natural resource administration, and emergency response. The projected value of the satellite-based Earth Observation market is anticipated to attain 11.3 billion USD by 2031 [1]. The Internet of Things (IoT) has resulted in substantial progress in information and communication technologies, hence augmenting the need for lucrative value-added services and applications. The proliferation of extensive Earth Observation data has been enabled by advancements in sensor, transmission, and storage technologies. This EOdataset has substantial information, although it poses challenges due to its unique attributes, including discrete observations and complex dimensions. The analysis of image retrieval methodologies seeks to tackle the difficulties presented by complex data [2]. The research highlights several challenges, including the limited availability and variety of training data, as well as the incapacity of existing algorithms to replicate human cognitive functions.
@@
Although there is a large amount of EO data available, there is a lack of effective conversion into actionable information for specialized sectors such as agriculture or disaster response. Big data systems that have frequent satellite revisits have the potential to revolutionize real-time monitoring [3]. NewSpace constellations [4] are expanding their capabilities beyond the provision of affordable and frequent data. They focus on addressing particular requirements by using specialized satellites and providing data analysis in addition to raw data, serving a broader spectrum of consumers. Furthermore, monitoring strategies need careful consideration of both geographical and temporal dimensions [5]. The extent of the phenomena being examined might vary from a single spot to an intricate three-dimensional grid. Resolution, which refers to the level of detail in measurements, and expanse, which indicates the region covered, are both important aspects. An example of researching the influence of topography on soil moisture may include the measurement of moisture levels at many locations along a slope, necessitating a high level of spatial resolution. Consequently, current EO systems are facing difficulties in managing the substantial volume of data being sent and fulfilling the diverse demands for the latency at which the data has to be processed. In order to address these significant challenges, the primary strategies involve two actions: firstly, improving the bandwidth of satellite communication networks that facilitate EO systems, and secondly, reducing data volume by employing on-board processors on the satellites to preprocess the EO data.
@@
Utilizing the inter-satellite links (ISLs) of the multi-satellite, multi-orbit connectivity systems has been seen as a potential strategy, as stated in the first direction [6]. The integration of several satellites and orbits, spearheaded by non-geostationary satellites (NGSO), is on the verge of transforming worldwide communication. Unlike conventional geostationary satellites (GSO), these systems provide reduced latency and loss as a result of their lower orbits. This research presented in [7] examines the communication in multi-satellite, multi-orbit connectivity systems, with a special emphasis on NGSO technology. It investigates the possibilities of this technology while recognizing the obstacles associated with its adoption. The data acquired by EO satellites may undergo processing, storage, and transmission to Earth using SatCom satellites positioned in various orbits. By using the ISLs, several routing channels may be constructed between the remote-sensing satellites and their respective ground stations. The technique of [8] may effectively handle the problem of overload and backhaul bottlenecks that occur at certain nodes in inter-satellite communication (ISC) networks, resulting in improved network capacity.
@@
In relation to the second aspect, the use of semantic communication (SemCom) principles has been recognized as a favorable strategy. The primary purpose of SemCom is to prioritize semantic accuracy above reducing bit/symbol error rates at the receiver side, which is in complete alignment with the core goals of most EO methods. The first principle of SemCom is the extraction of semantic data, wherein pertinent information is initially obtained, condensed, or broken down from unprocessed data. Afterward, just the information related to meaning is transferred to the central core. The PhiSat-1 project, presented in [12] by the European Space Agency (ESA), was among the first demonstrations of this concept. With the recent launch of PhiSat-2, this approach is now advancing to new capabilities. Two 6U CubeSats are deployed in this mission to capture Earth photographs. The data is processed inside the CubeSats to filter out information depending on the cloud coverage of the scene. The second concept revolves around semantic-oriented network management, specifically SemCom. SemCom’s objective is to oversee all network functionalities across various layers (PHY, MAC, Network, Transportation, Application) by taking into account the semantics of the data and the requirements of applications related to EO tasks [13].
Figure 1: The system model of semantic-oriented EO satellite networks.
@@
Our approach introduces a systematic method for optimizing data flow and processing across multi-orbit EO networks, offering solutions to the major challenges of managing large-scale, diverse EO datasets. The purpose of presenting Fig. 1 is to visually illustrate the architecture of the proposed multi-orbit inter-satellite network and to clarify the layered structure of satellite communication and data processing. This architecture helps contextualize our approach by demonstrating how Earth observation satellites in various orbits (LEO, MEO, GEO) can coordinate to gather and integrate data from ground sensors in diverse environments, such as agricultural fields and maritime regions. By providing this visual model, we aim to make the complex network of multi-orbit data flow and processing more accessible, highlighting the systematic approach we propose for optimizing data integration within Earth observation systems. The EO data, initially processed for semantic content from its raw form, is then transmitted to a ”Ground segment”, which includes data centers and user control interfaces for further analysis and user access. Data travels through multiple layers of communication within the network. At the base level, the PHY Layer (Physical Layer) ensures the fundamental transmission of data, followed by the MAC Layer (Medium Access Control) responsible for resource allocation. Moving upwards, the Network Layer manages the interconnectivity across the satellite network. The Transportation Layer handles the movement of data across the system, connecting the ground segments to various orbital satellites.
@@
Above the operational layers is the Semantic Effectiveness Plane, which oversees and analyzes the flow of semantic data. The model features two key processes: Semantic Extraction and Semantic Restoration. At the ground level, data extracted from sensors undergoes semantic processing—either through ground-based AI servers or onboard edge computing—where meaningful patterns and insights are identified. This processed semantic information is subsequently restored into actionable applications, ranging from environmental monitoring to disaster management. The model proposes a closed-loop system in which raw data from ground sensors is transformed into semantically rich insights, facilitating seamless communication between satellite networks in space and ground infrastructure.
We summarize our contributions as follows.
  • We present an in-depth review of the latest developments in deep learning for semantic data processing in EO systems and 6G communication, emphasizing on-board processing and integration for segmentation, data compression, and pixel-level categorization to improve remote sensing analysis and retrieval. Additionally, we examine the evolution of SemCom, highlighting efficiency improvements through joint source-channel coding (JSCC) and deep learning, while exploring frameworks that enhance LLM adaptability and decision-making in networking and EO using task-specific models and advanced data fusion.
  • The cognitive semantic EO system model enhances satellite networks by leveraging Discrete Task-Oriented JSCC (DT-JSCC) [14] and Semantic Data Augmentation (SA) [15] techniques to transmit task-relevant semantic features rather than raw data, significantly reducing communication overhead and energy consumption in 6G satellite communication.
  • The proposed framework incorporates inter-satellite communication and deep learning models for on-board processing, optimizing the transmission, analysis, and decision-making of multi-spectral EO data through an end-to-end semantic-augmented communication process, leading to improved classification accuracy and system efficiency.
The rest of the paper is structured as follows. Section II discusses semantic data processing from semantic data compression and fusion points of view to enhance the performance of applications that are based on EO data. Section III presents the evolution of SemCom networks, their attributes, and source and channel coding aspects with a framework for semantic inference-based EO. Section IV discusses the role of LLMs in enhancing domain-specific semantic effectiveness and automating EO. Section V introduces the SemCom-based, resource-efficient EO system, showcasing advancements in on-board processing hardware—such as SRAM-based FPGAs, MPSoCs, and reconfigurable designs—that pave the way for adaptive, resilient, and energy-efficient satellite systems capable of meeting the dynamic demands of modern space missions. Section VI concludes the work with a future outlook.
II Semantic data processing for EO applications
This section delves into the concept of semantic segmentation, a technique that assigns precise identifiers to individual pixels in photographs. It concentrates on the application of this methodology to the examination of satellite imagery and other types of EO data. The content explores the application of semantic segmentation in a variety of applications, including trajectory facilitation, information retrieval, and picture compression. We also underscore the current advancements in deep learning for the classification of various EO data types using semantic knowledge. Irrespective of the intended meaning of the source information, traditional communication, based on the Shannon theory, emphasizes the reliable reception of individual sent bits. In contrast, the SemCom focuses on the successful transmission of only necessary semantic information to convey the meaning of the source message and thus it aspires to achieve the information capacity beyond the traditional Shannon paradigm [16]. Semantic-information feature learning [17] is the key to addressing the utilization of cognitive techniques employed as a means to direct computational resources towards activities of higher importance. The process of achieving dynamic adaptation in semantic compression involves the utilization of a feature learning module and an attention feature module. These modules enable the source encoder to generate a variable number of symbols and modify the capabilities of the source encoder and channel encoder through the use of cognitive techniques. The evolution of semantic communication [18] is traced back to the early 20th century with its continuous growth into the realm of modern communications beyond 5G and 6G technologies. In light of this, there is a significant need to develop more intelligent and efficient communication protocols that can meet the diverse quality of service (QoS) needs. This must be done while addressing the challenge of limited communication bandwidth and computation. The development of an intelligent communication system is considered essential in both industry and academia. Such a system is not limited to memorizing data flows based on rigorous regulations but also aims to comprehend, analyze, and articulate the fundamental semantics. In order to facilitate the implementation of immediate intelligent services in future communication systems, it is advantageous to distill and communicate only the information that is pertinent to the task at hand and precise semantics. The semantic approach effectively reduces the overall latency of the system.
@@ II Semantic data processing for EO applications
Achieving accurate categorization at the pixel level is crucial in applications like driverless cars, where the capacity to understand lane markings and precisely identify the position of objects is essential. However, the process of semantic segmentation may be complicated by barriers such as hidden objects or congested backgrounds. Notwithstanding these challenges, semantic segmentation offers benefits that go beyond autonomous vehicles. The technology enables the examination of medical images by identifying abnormalities, assists robots in navigating their surroundings, and improves video surveillance by detecting potentially troublesome behavior. Recent progress in deep learning has led to remarkable achievements in achieving high precision at the pixel level [19]. However, its performance is inferior to that of normal satellite images when dealing with non-standard data such as point clouds and spectrum imaging. Likewise, a significant disparity in performance occurs when dealing with limited data sets. The lack of comprehensively classified, unconventional remote sensing data hinders the development and evaluation of innovative deep-learning methods.
@@
II-1 Computational semantic compression
The study presented in [20] investigates potential resolutions for constraints such as streamlining features, indexing, semantic labeling, and ultimately, approaches for formulating and executing searches. As stated in [21], the primary objective of the study is to evaluate the existing condition of research on extracting information from EO data. The authors of [22] provide a new notion of a semantic EO data cube. It enriches traditional exploratory optimization data cubes by including an extra layer, referred to as a ”nominal” or categorical interpretation, for each observation. Thus, it becomes feasible to use semantic searches that directly relate to the substance of the data and are easily comprehensible to people. By using machine-based reasoning, this approach improves automated analysis and information extraction in comparison to non-semantic cubes. Fundamentally, semantic EO data cubes enable a wider range of users to utilize comprehensive EO data. This study assesses CORSA [23] and [34], a novel compression technique driven by artificial intelligence that produces valuable data representations. The authors contrast conventional water detection methods with innovative techniques that use the compressed data and representations generated by CORSA. The findings demonstrate that CORSA compression does not negatively impact accuracy and enables the direct establishment of water detection networks using its representations. The versatility of CORSA provides a viable option for the processing and compression of data onboard, tackling the growing difficulty of handling Earth Observation (EO) data. Nonetheless, compression is intrinsically application-specific, which poses challenges when striving to create diverse apps that share a similar dataset. In this context, the authors of [24] suggest an innovative approach using semantic technology to establish a standardized language for Earth Observation data. This technique improves the integration of raster imagery data, such as land cover maps, with other geographical information, hence enhancing the efficiency of analysis and the use of this essential EO data.
II-2 Fusion of semantic trajectories
Global Positioning System (GPS) data produces huge amounts of trajectory information. While scholars have largely focused on the preservation and analysis of these trajectories, there is a growing need to understand the underlying importance of the movement. The material includes traditional compression algorithms, the criteria for defining moving objects, approaches for measuring compression efficiency, and potential future applications. For real-time applications, conventional methods for producing substantial movement data are inadequately fast. The research presented in [26] discusses a novel approach that combines mobility data with geography and application information to produce semantic trajectories. These improved trajectories not only reveal the destinations of objects but also provide a valuable understanding of the fundamental causes of their motion.
II-3 Advancements in Semantic Comprehension